CONVERT FROM TRANSFER FUNCTION FORM TO STATE SPACE FORM AND COMPUTE THE EIGENVALUES AND EIGENVECTORS

Y(s) U(s) ー0.0 1785G3 +77.88 +0.496s +0.446) (s2 +0.00466s +0.0053)(s2+0.806s+1.311)
Answer #1

we do as following

y(s)[s4 +0.81066+1.3200596s2 +0.0069483]=u(s)[-0.01785s3 -1.38873s2 -0.0088536s-0.0079611]

this gives

s4y(s)+1.3200596s2 y(s)+0.0069483y(s)+0.81066y(s)=-0.01785s3 u(s)- 1.38873s2 u(s) -0.0088536su(s)-0.0079611u(s)

this gives the following differential equation

y+0.81066y+1.32005596y+0.0069483y=-0.01785u-1.38873u-0.0088536u-0.0079611